OpenStudy (anonymous):

The average number of red, blue, and green beanbags in a store is 136. There are 30 more red beanbags than blue beanbags. There are 15 fewer green beanbags than blue beanbags. How many green beanbags are in the store? Use bar models to help you

OpenStudy (sgadi):

red = 70.3333 blue = 40.3333 green = 25.3333

OpenStudy (anonymous):

how did u come up with this answer?

OpenStudy (sgadi):

r+b+g=136 r=b+30 g=b-15 by solving above 3 equations ...

OpenStudy (anonymous):

are you allowed to have decimals in your answer though? because a beanbag is a whole thing...

OpenStudy (anonymous):

probably not sons 4th grade math class

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

sgadi answer is what i came up with but son said wasn't sure if decimals were allowed.

OpenStudy (anonymous):

yeah...im the same

OpenStudy (sgadi):

then ... I guess ... there is no solution for this problem